a company just paid an annual dividend of $1.20 per share. the share is priced at 24$ and the required rate of return is 11.02%. what is the dividend growth rate ?
Answer #1

Required return=(D1/Current price)+Growth rate

0.1102=(1.2*(1+Growth rate)/24+Growth rate

0.1102=(1.2+1.2*Growth rate)/24+Growth rate

0.1102=0.05+0.05*Growth rate+Growth rate

Growth rate=(0.1102-0.05)/(0.05+1)

=5.73%(Approx)